17. A business worth $180,000 is expected to grow at 12% per year compounded annually for the next 4 years. (a) Find the expected future value. (b) If funds from the sale of the business today would be placed in an account yielding 8% compounded semiannually, what would be the minimum acceptable price for the business at this time?

18. A corporation worth $40 million is expected to grow at 8% per year compounded annually for 5 years. (a) Find the future value to the nearest million. (b) The owners then propose to sell the firm and invest the proceeds in a new venture that should grow at 12% compounded annually for 4 years. Beginning with the future value from part (a) rounded to the nearest million, find the expected future value to the nearest million at the end of 4 additional years.
Answer #1

Present value = $180000

Growth rate = 12%

Time = 4 years

FV=PV (1+r)^n

= 180000 (1 + 0.12)^4

= 180000 (1.12)^4

=180000 (1.575)

Expected future value = $ 283233.5